Circular No. 7481

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000da IN UGC 5 A. Pastorello, E. Giro, G. Altavilla, and S. Benetti, Padua Astronom-

ical Observatory, report that inspection of a partially reduced, noisy CCD spectrogram (range 325–752 nm; resolution 1.7 nm) obtained with the Asiago 1.8-m telescope (+ Afosc), on Aug. 23.03 UT reveals that SN 2000da (cf. IAUC 7474) is a type-II supernova caught soon after maximum light. The spectrum is dominated by a blue continuum and a broad and slightly asymmetric Hα emission (FWHM ∼ 9500 km/s). Hβ and Fe II lines show instead well-developed P-Cyg profiles, and an expansion velocity of ∼ 7800 km/s is derived from the minimum of Hβ absorption.

SUPERNOVA 2000db IN NGC 3949 Pastorello et al. also report that preliminary inspection of a partially

reduced CCD spectrogram (range 325–752 nm; resolution 1.7 nm), obtained as above on Aug. 22.82 UT, shows that SN 2000db (cf. IAUC 7475) is a type-II supernova, several weeks after outburst. Strong P-Cyg lines of H, Fe II (multiplets 40 and 42), Sc II, Na I (D), and Ba II are superimposed on a relatively red continuum. The expansion velocity deduced from the minimum of Hα line is ∼ 7300 km/s.

K. Hornoch, Lelekovice, Czech Republic, reports that his CCD Rband photometry (0.35-m reflector; comparison stars GSC 3452.1250 and 3452.1219) of SN 2000db shows that its brightness decreased linearly by ∼ 0.013 mag/day between Aug. 14.84 and 20.82 UT.

SUPERNOVA 2000de IN NGC 4384 Altavilla, Pastorello, Benetti, M. Turatto, and Giro also report that

the inspection of a preliminarily reduced spectrum (range 325–750 nm; resolution 1.7 nm) taken by C. Tappert with the Asiago 1.8-m telescope (+ Afosc) on Aug. 23.85 UT, indicates that SN 2000de (cf. IAUC 7478) is a type-Ib supernova, 2–3 weeks after maximum light. The spectrum shows a red continuum and the following strong P-Cyg lines of He I: 587.6 (dominant), 667.8, and 706.5 nm. The average expansion velocity deduced from the minima of He I lines is ∼ 7500 km/s.

COMET C/2000 N2 (SOHO) Addendum. On IAUC 7459, it should be added that P. Shkreby also

independently found C/2000 N2 in real-time SOHO website data.

2000 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green